Qd= 20 - 2P



Qs= 10 + 3P




a. Determine the equilibrium price and quantity.



b. Calculate the point price elasticity of demand at equilibrium



C. Calculate the point price elasticity of supply at equilibrium



d. Suppose an indirect tax of $10 per unit is imposed on this commodity, find the now equilibrium price and quantity of the product.




1
Expert's answer
2022-05-31T10:04:23-0400

"Q_d=20-2P,\\
Q_s=10+3P,"

"Q_d=Q_s,"

"P=2,"

"Q_d=Q_s=16."